OpenStudy (keeli_grace):

Which is a correct interpretation of the slope of the line in this graph? A. The volume of water in the tank is decreasing by 2.5 L per hour. B. The volume of water in the tank is decreasing by 0.2 L per hour. C. The volume of water in the tank is decreasing by 0.4 L per hour. D. The volume of water in the tank is decreasing by 2 L per hour.

OpenStudy (keeli_grace):

Here is the graph:

OpenStudy (keeli_grace):

@Gabylovesyou @Burningitdown

OpenStudy (ranga):

Slope of the line = \(\Large \frac{y_2-y_1}{x_2-x_1}\).

OpenStudy (ranga):

\(\Large x_1 = 2, y_1=8; x_2 =7, y_2 = 6\)

OpenStudy (ranga):

Since the slope of the graph is negative, the volume of water in the tank decreases with time. The above formula will tell you by how much it decreases each hour.

OpenStudy (keeli_grace):

I don't understand @ranga

OpenStudy (ranga):

You are given two points on the straight line: (2,8) and (7,6). From these two points you can calculate the slope of the graph. The formula is given above. Just plug in the numbers and calculate the slope. That will be the rate at which the volume of water in the tank is decreasing every hour.

OpenStudy (keeli_grace):

But the answer is B. I can tell not by figuring out the problem, but instead by implying the picture the answer choices.

OpenStudy (ranga):

I don't think B is the correct answer. Slope = \(\Large \frac{6-8}{7-2} = \frac {-2}{5} = -0.4\). The negative sign tell us the volume is decreasing. And it is decreasing at the rate of 0.4 L per hour.

OpenStudy (keeli_grace):

The system said I got it right though...?

OpenStudy (ranga):

Calculations shown above indicates the correct answer is C. Sometimes there are errors in the system.